OpenStudy (anonymous):

I need help understanding how to solve a linear equation with fractions 1/6 (2/c -1/6) = 1/3c -1/6

OpenStudy (turingtest):

is the problem\[\frac16(\frac2c-\frac16)=\frac1{3c}-\frac16\]?

OpenStudy (anonymous):

yes

OpenStudy (turingtest):

you need to find the GCF of the denominators and multiply the whole equation by that to get rid of the fractions do you know how to find the GCF of a set of expressions?

OpenStudy (anonymous):

that would be 6?

OpenStudy (turingtest):

actually it would be 6c if we don't include the c, then we leave out c as a factor and we will still have fractions, which we don't want

OpenStudy (turingtest):

the denominators are 6=2*3 c 3 so their GCF is 2*3*c=6c

OpenStudy (turingtest):

now multiply everything by 6c

OpenStudy (anonymous):

okay could you show me how to set the problem up to be solved

OpenStudy (turingtest):

\[6c[\frac16(\frac2c-\frac16)]=6c[\frac1{3c}-\frac16]\]now simplify and solve for c

OpenStudy (anonymous):

so i can cancel out the 6 with 1 and cancel out the 6c leaves 1c, then i multiplu 1c x 2 =2/c

OpenStudy (anonymous):

couls you show what it look like after i simplify plz

OpenStudy (turingtest):

what is\[6c(\frac16)\]?

OpenStudy (anonymous):

cancel out the bottom 6 by 6 and get 1, cancel out the 6c with 6 and get 1c

OpenStudy (anonymous):

1/6 (2c-1) =1/3c-1/6 this is how it looks. Srry i made a mistake with the equation